# GeoClip: Geometry-Aware Clipping for Differentially Private SGD

### v1

@@ 1 Introduction
As machine learning models are increasingly trained on sensitive user data, ensuring strong privacy guarantees during training is essential to reduce the risk of misuse, discrimination, or unintended data exposure. Differential privacy (DP) (Dwork et al., 2006a, b) offers a principled framework for protecting individual data, and has become a cornerstone of privacy-preserving machine learning. In deep learning, the most widely used approach for DP is differentially private stochastic gradient descent (DP-SGD) (Abadi et al., 2016), which clips per-sample gradients and adds calibrated noise to their average.
Despite its widespread use, standard DP-SGD has a key limitation: it relies on a fixed clipping threshold to bound the sensitivity of individual gradients. Selecting this threshold poses a challenging privacy-utility tradeoff—setting it too low discards useful gradient information, while setting it too high increases sensitivity and necessitates injecting more noise, ultimately degrading model performance. This trade-off was observed empirically by McMahan et al. (2018) and later analyzed theoretically and shown to be a fundamental limitation of differentially private learning by Amin et al. (2019). Moreover, the optimal threshold can vary over the course of training, across tasks, and between datasets, limiting the effectiveness of a fixed setting.
To address this, recent work has proposed adaptive strategies that dynamically adjust the clipping threshold during training. One class of methods uses decay schedules to reduce the threshold over time. Yu et al. (2018) and Du et al. (2021) propose linear and near-linear decay rules, respectively, where the schedule is predefined and does not depend on the dataset—hence, no privacy budget is required. Lin et al. (2022) introduce a nonlinear decay schedule, along with a transfer strategy that leverages public data to guide threshold selection. Recently, methods have been introduced which set the clipping level based on the data during the training process. These include AdaClip (Pichapati et al., 2019), which applies coordinate-wise clipping based on estimated gradient variances. Yet another is quantile-based clipping (Andrew et al., 2021), which sets the threshold using differentially private quantiles of per-sample gradient norms. Although (Andrew et al., 2021) was designed for federated learning, it can be adapted to centralized DP-SGD. These adaptive strategies have been shown to improve model utility while preserving privacy guarantees.
@@
A different approach to correlations in gradients from the literature considers introducing correlations—dependencies across iterations and between entries of the noise vector—into the noise, rather than injecting i.i.d. Gaussian noise. This has been shown to improve the utility of private training (Denisov et al., 2022; Kairouz et al., 2021)Choquette-Choo et al. (2024) strengthen this direction by analytically characterizing near-optimal spatio-temporal correlation structures that lead to provably tighter privacy-utility tradeoffs. However, these approaches have been developed independently of adaptive clipping, and the effect of combining both methods remains unexplored. In addition, (Choquette-Choo et al., 2024) uses pre-determined correlations in the noise, rather than being tailored to the data, as our approach is.
@@ 2 General Framework
GeoClip generalizes AdaClip of Pichapati et al. (2019), which itself extends standard DP-SGD. AdaClip essentially assumes that $M_{t}$ is diagonal for per-coordinate scaling, whereas GeoClip allows $M_{t}$ to be any full-rank matrix. This added flexibility allows GeoClip to account for correlations between gradient components and inject noise along more meaningful directions.
A key feature of our framework is that the privacy guarantees remain unaffected by the choice of $M_{t}$ and $a_{t}$, as long as no privacy budget is used to compute them. This allows $M_{t}$ and $a_{t}$ to be chosen entirely based on utility, without compromising privacy. The main challenge, then, is how to select these parameters effectively. To address this, we first define a performance metric for GeoClip. Inspired by Pichapati et al. (2019), we derive the following convergence bound for our framework and use it to guide the design of $M_{t}$ and $a_{t}$, ultimately improving both convergence and training efficiency.
@@
The above result generalizes Theorem 2 in Pichapati et al. (2019); proof details are in the supplement.
@@
The proof of Theorem 2 is in the supplement.
@@ 3 Algorithm Overview
We then perform an SVD on $Z$ and retain the top $k$ singular vectors and squared singular values as the updated eigenvectors and eigenvalues. We note that this computation takes $\mathcal{O}(dk^{2}+k^{3})$ time Li et al. (2019), highlighting that it is only linear in $d$. The rest of the procedure follows Algorithm 1 by replacing the moving average over the full covariance matrix in line 12 of Algorithm 1 with the low-rank approximation described above. Another necessary change to the algorithm is the following: since $U_{t}$ is no longer a square matrix, the transformation $M_{t}$ takes the gradient into the lower $k$-dimensional space to clip and add noise, such that the resulting $M_{t}^{\text{inv}}$ returns to the full $d$-dimensional space. The complete version of this variant is provided as Algorithm 3 in the supplementary material.
@@ 4 Experimental Results
We present empirical results demonstrating the benefits of our proposed GeoClip framework compared to AdaClip, quantile-based clipping, and standard DP-SGD. For quantile-based clipping, we use the median of per-sample norms, which has been shown to perform well across various learning tasks Andrew et al. (2021). All $(\varepsilon,\delta)$-DP guarantees are computed using the Connect-the-Dots accountant Doroshenko et al. (2022). All experiments were conducted on Google Colab using CPU resources.
@@
Our results indicate that our framework in Algorithm 1 is robust to the choice of hyperparameters $\beta_{1}$, $\beta_{2}$, and $h_{1}$. Standard values commonly used in optimization, such as $\beta_{1}=0.99$ and $\beta_{2}=0.999$, work well in our setting. The parameter $h_{1}$ only needs to be a small positive constant (e.g., $10^{-15}$) to ensure numerical stability. Since the eigenvalues are clamped to the range $[h_{1},h_{2}]$ (line 14 in  Algorithm 1), the trace term $\sum_{i}\sqrt{\lambda_{i}}$ is bounded between $d\sqrt{h_{1}}$ and $d\sqrt{h_{2}}$, where $d$ is the dimensionality. This allows us to set the parameter $\gamma$ to $1$, as its effect can be absorbed by tuning $h_{2}$. For $h_{2}$, we have observed that the values $1$ and $10$ perform consistently well across datasets. Throughout all experiments, we tune only $h_{2}$ to select between these two options. A similar setup is used in Algorithm 2 with $\beta_{2}$ replaced by $\beta_{3}$ (set to $0.99$), which is also robust across experiments.
@@ 4.2 Tabular Datasets
In addition to the synthetic dataset, we run experiments to compare the performance of different clipping strategies on three real-world datasets: Diabetes Efron et al. (2004), Breast Cancer Wolberg et al. (1993), and Android Malware Borah and Bhattacharyya (2020). We briefly describe each dataset and its corresponding learning task below.
@@
Table 1: Diabetes dataset test MSE comparison for $\delta=10^{-5}$, batch size $=32$, and model dimension $d=11$.
@@
GeoClip (ours) 0.073$\pm$ 0.015 0.044$\pm$0.003 0.039$\pm$0.009
AdaClip 0.077$\pm$0.027 0.062$\pm$0.028 0.055$\pm$0.014
Quantile 0.090$\pm$0.027 0.083$\pm$0.044 0.072$\pm$0.014
DP-SGD 0.108$\pm$0.040 0.095$\pm$0.047 0.072$\pm$0.040
Table 2: Breast Cancer dataset test accuracy (%) comparison for $\delta=10^{-5}$, batch size $=64$, and model dimension $d=62$.
@@
GeoClip (ours) 87.87$\pm$3.32 88.57$\pm$3.37 93.63$\pm$1.63
AdaClip 84.90$\pm$5.91 85.42$\pm$5.34 87.71$\pm$5.87
Quantile 81.41$\pm$12.53 81.63$\pm$10.71 92.28$\pm$2.68
DP-SGD 77.32$\pm$6.17 79.42$\pm$9.71 85.95$\pm$4.45
Table 3: Malware dataset test accuracy (%) comparison for $\delta=10^{-5}$, batch size $=512$, and model dimension $d=484$.
@@ 4.3 Final Layer Fine-Tuning
To demonstrate this, we design an experiment where a convolutional neural network (CNN) is first trained on MNIST LeCun et al. (1998) using the Adam optimizer and then transferred to Fashion-MNIST Xiao et al. (2017) by freezing all layers except the final fully connected layer. The CNN consists of two convolutional and pooling layers followed by a linear compression layer that reduces the feature size to 50, resulting in a total of only 510 trainable parameters. We fine-tune this layer using different methods under varying privacy budgets, and present the results in Table 4.
@@ 4.4 Low-Rank PCA Results
We also evaluate our method on the USPS dataset Hull (1994) using logistic regression with 2,570 trainable parameters (256 input features $\times$ 10 classes + 10 biases). The USPS dataset contains 9,298 grayscale handwritten digit images (0–9), each of size $16\times 16$ pixels. This compact benchmark is commonly used for evaluating digit classification models. For this dataset, we apply a low-rank PCA approximation with rank 100. Results are shown in the right panel of Figure 2. As with the synthetic dataset, GeoClip with low-rank PCA also achieves faster convergence on USPS compared to baseline methods. We provide additional $\varepsilon$-vs-iteration plots for both datasets, illustrating how faster convergence reduces overall privacy cost in the supplementary material.
[…]
@@ 5 Conclusions
Limitations.
The linear transformation involved in GeoClip requires an additional computation via an eigendecomposition. Our low-rank approximation addresses that to some extent. The algorithm introduces additional hyperparameters (most notably $h_{2}$) compared to standard DP-SGD, which must be tuned for optimal performance. Our experiments have been performed on a limited collection of datasets; additional testing is needed to see how our algorithm performs in more generality.
Broader impact.
As discussed in the Introduction, our work is motivated by societal concerns, with a focus on improving the theoretical limits of differentially private optimization.
@@ Appendix A Proof of Theorem 1
Summing over $t=0$ to $T-1$ and applying the telescoping sum, we get:
@@
$+\eta\sum_{t=0}^{T-1}\ \mathbb{E}\left[\beta(a_{t})\left(\operatorname{Tr}\left(\text{Cov}\left(M_{t}g_{t}\middle|\theta^{t}\right)\right)+\|M_{t}(\mathbb{E}[g_{t}|\theta^{t}]-a_{t})\|^{2}\right)\right]$ (61)
@@ Appendix C GeoClip with Low-Rank PCA
[…]

### v2

@@ 1 Introduction
As machine learning models are increasingly trained on sensitive user data, ensuring strong privacy guarantees during training is essential to reduce the risk of misuse, discrimination, or unintended data exposure. Differential privacy (DP) [1, 2] offers a principled framework for protecting individual data, and has become a cornerstone of privacy-preserving machine learning. In deep learning, the most widely used approach for DP is differentially private stochastic gradient descent (DP-SGD) [3], which clips per-sample gradients and adds calibrated noise to their average.
Despite its widespread use, standard DP-SGD has a key limitation: it relies on a fixed clipping threshold to bound the sensitivity of individual gradients. Selecting this threshold poses a challenging privacy-utility tradeoff—setting it too low discards useful gradient information, while setting it too high increases sensitivity and necessitates injecting more noise, ultimately degrading model performance. This trade-off was observed empirically by McMahan et al. [4] and later analyzed theoretically and shown to be a fundamental limitation of differentially private learning by Amin et al. [5]. Moreover, the optimal threshold can vary over the course of training, across tasks, and between datasets, limiting the effectiveness of a fixed setting.
To address this, recent work has proposed adaptive strategies that dynamically adjust the clipping threshold during training. One class of methods uses decay schedules to reduce the threshold over time. Yu et al. [6] and Du et al. [7] propose linear and near-linear decay rules, respectively, where the schedule is predefined and does not depend on the dataset—hence, no privacy budget is required. Lin et al. [8] introduce a nonlinear decay schedule, along with a transfer strategy that leverages public data to guide threshold selection. Recently, methods have been introduced which set the clipping level based on the data during the training process. These include AdaClip [9], which applies coordinate-wise clipping based on estimated gradient variances. Yet another is quantile-based clipping [10], which sets the threshold using differentially private quantiles of per-sample gradient norms. Although [10] was designed for federated learning, it can be adapted to centralized DP-SGD. These adaptive strategies have been shown to improve model utility while preserving privacy guarantees.
@@
A different approach to correlations in gradients from the literature considers introducing correlations—dependencies across iterations and between entries of the noise vector—into the noise, rather than injecting i.i.d. Gaussian noise. This has been shown to improve the utility of private training [11, 12]Choquette-Choo et al. [13] strengthen this direction by analytically characterizing near-optimal spatio-temporal correlation structures that lead to provably tighter privacy-utility tradeoffs. However, these approaches have been developed independently of adaptive clipping, and the effect of combining both methods remains unexplored. In addition, [13] uses pre-determined correlations in the noise, rather than being tailored to the data, as our approach is.
@@ 2 General Framework
GeoClip generalizes AdaClip of Pichapati et al. [9], which itself extends standard DP-SGD. AdaClip essentially assumes that $M_{t}$ is diagonal for per-coordinate scaling, whereas GeoClip allows $M_{t}$ to be any full-rank matrix. This added flexibility allows GeoClip to account for correlations between gradient components and inject noise along more meaningful directions.
A key feature of our framework is that the privacy guarantees remain unaffected by the choice of $M_{t}$ and $a_{t}$, as long as no privacy budget is used to compute them. This allows $M_{t}$ and $a_{t}$ to be chosen entirely based on utility, without compromising privacy. The main challenge, then, is how to select these parameters effectively. To address this, we first define a performance metric for GeoClip. Inspired by Pichapati et al. [9], we derive the following convergence bound for our framework and use it to guide the design of $M_{t}$ and $a_{t}$, ultimately improving both convergence and training efficiency.
@@
The above result generalizes Theorem 2 in Pichapati et al. [9]; proof details are in the supplement.
@@
The proof of Theorem˜2 is in the supplement.
@@ 3 Algorithm Overview
We then perform an SVD on $Z$ and retain the top $k$ singular vectors and squared singular values as the updated eigenvectors and eigenvalues. We note that this computation takes $\mathcal{O}(dk^{2}+k^{3})$ time [14], highlighting that it is only linear in $d$. The rest of the procedure follows Algorithm 1 by replacing the moving average over the full covariance matrix in line 12 of Algorithm˜1 with the low-rank approximation described above. Another necessary change to the algorithm is the following: since $U_{t}$ is no longer a square matrix, the transformation $M_{t}$ takes the gradient into the lower $k$-dimensional space to clip and add noise, such that the resulting $M_{t}^{\text{inv}}$ returns to the full $d$-dimensional space. The complete version of this variant is provided as Algorithm 3 in the supplementary material.
@@ 4 Experimental Results
We present empirical results demonstrating the benefits of our proposed GeoClip framework compared to AdaClip, quantile-based clipping, and standard DP-SGD. For quantile-based clipping, we use the median of per-sample norms, which has been shown to perform well across various learning tasks [10]. All $(\varepsilon,\delta)$-DP guarantees are computed using the Connect-the-Dots accountant [15]. All experiments were conducted on Google Colab using CPU resources.
@@
Our results indicate that our framework in Algorithm˜1 is robust to the choice of hyperparameters $\beta_{1}$, $\beta_{2}$, and $h_{1}$. Standard values commonly used in optimization, such as $\beta_{1}=0.99$ and $\beta_{2}=0.999$, work well in our setting. The parameter $h_{1}$ only needs to be a small positive constant (e.g., $10^{-15}$) to ensure numerical stability. Since the eigenvalues are clamped to the range $h_{1},h_{2}]$ (line 14 in  Algorithm˜1), the trace term $\sum_{i}\sqrt{\lambda_{i}}$ is bounded between $d\sqrt{h_{1}}$ and $d\sqrt{h_{2}}$, where $d$ is the dimensionality. This allows us to set the parameter $\gamma$ to $1$, as its effect can be absorbed by tuning $h_{2}$. For $h_{2}$, we have observed that the values $1$ and $10$ perform consistently well across datasets. Throughout all experiments, we tune only $h_{2}$ to select between these two options. A similar setup is used in Algorithm˜2 with $\beta_{2}$ replaced by $\beta_{3}$ (set to $0.99$), which is also robust across experiments.
@@ 4.2 Tabular Datasets
In addition to the synthetic dataset, we run experiments to compare the performance of different clipping strategies on three real-world datasets: Diabetes [16], Breast Cancer [17], and Android Malware [18]. We briefly describe each dataset and its corresponding learning task below.
@@
Table 1: Diabetes dataset test MSE comparison for $\delta=10^{-5}$, batch size =32=32, and model dimension $d=11$.
@@
[…]
Table 2: Breast Cancer dataset test accuracy (%) comparison for $\delta=10^{-5}$, batch size =64=64, and model dimension $d=62$.
@@
[…]
Table 3: Malware dataset test accuracy (%) comparison for $\delta=10^{-5}$, batch size =512=512, and model dimension $d=484$.
@@ 4.3 Final Layer Fine-Tuning
To demonstrate this, we design an experiment where a convolutional neural network (CNN) is first trained on MNIST [19] using the Adam optimizer and then transferred to Fashion-MNIST [20] by freezing all layers except the final fully connected layer. The CNN consists of two convolutional and pooling layers followed by a linear compression layer that reduces the feature size to 50, resulting in a total of only 510 trainable parameters. We fine-tune this layer using different methods under varying privacy budgets, and present the results in Table 4.
@@ 4.4 Low-Rank PCA Results
We also evaluate our method on the USPS dataset [21] using logistic regression with 2,570 trainable parameters (256 input features $\times$ 10 classes + 10 biases). The USPS dataset contains 9,298 grayscale handwritten digit images (0–9), each of size $16\times 16$ pixels. This compact benchmark is commonly used for evaluating digit classification models. For this dataset, we apply a low-rank PCA approximation with rank 100. Results are shown in the right panel of Figure 2. As with the synthetic dataset, GeoClip with low-rank PCA also achieves faster convergence on USPS compared to baseline methods. We provide additional $\varepsilon$-vs-iteration plots for both datasets, illustrating how faster convergence reduces overall privacy cost in the supplementary material.
Figure 2: The left panel shows results on the synthetic Gaussian dataset with 400 features using a rank-50 PCA approximation for GeoClip. The plot displays average test accuracy $(\%)$ over 80 iterations with a batch size of 1024. GeoClip achieves the fastest convergence and highest average accuracy. The right panel shows results on the USPS dataset using a rank-100 approximation over 25 iterations with a batch size of 1024, where a similar convergence trend is observed. Shaded regions represent standard deviation across 20 random seeds.
@@ Appendix A Proof of Theorem 1
Summing over $t=0$ to T1T-1 and applying the telescoping sum, we get:
@@
$\hskip-11.38109pt+\eta\sum_{t=0}^{T-1}\ \mathbb{E}\left[\beta(a_{t})\left(\operatorname{Tr}\left(\text{Cov}\left(M_{t}g_{t}\middle|\theta^{t}\right)\right)+\|M_{t}(\mathbb{E}[g_{t}|\theta^{t}]-a_{t})\|^{2}\right)\right]$ (61)
@@ Appendix C GeoClip with Low-Rank PCA
Figure 3: The left panel shows results on the synthetic Gaussian dataset with 400 features using a rank-50 PCA approximation for GeoClip. The plot displays average test accuracy $(\%)$ over 80 iterations with a batch size of 1024. GeoClip achieves the fastest convergence and highest average accuracy. The right panel shows results on the USPS dataset using a rank-100 approximation over 25 iterations with a batch size of 1024, where a similar convergence trend is observed. Shaded regions represent standard deviation across 20 random seeds.